# LLMAID: Identifying AI Capabilities in Android Apps with LLMs

### v1

@@ 3. Evaluation
In this section, we evaluate the effectiveness of our proposed approach, LLMAID, by addressing the following research questions:
@@
    This research question aims to assess the accuracy and reliability of LLMAID in analyzing AI capabilities within previously recognized AI Android apps. Addressing this question validates the fundamental effectiveness of LLMAID and provides the foundation for subsequent evaluations.
@@
    We explore the clarity and informativeness of the AI service summaries generated by our framework. These summaries are intended to provide end users, developers, and regulators with an interpretable overview of the AI functionalities embedded within a given Android application.
@@ 3.1. Dataset
We perform our experiments on two separate datasets. The first comes from the existing literature, which was collected from AndroZoo (allix2016androzoo) and contains 56,682 AI Apps determined by a rule-based approach. The other is self-curated, comprising 15,517 Android apps, and reflects an open-world setting to address the absence of newer released applications, particularly those potentially integrating with emerging LLMs such as ChatGPT. Detailed descriptions of these two datasets are provided in the following paragraphs.
3.1.1. Dataset 1
@@
Li et al. (li2022ai) proposed a SOTA rule-based approach, AI Discriminator, to identify AI-driven Android apps via keywords matching on the dissected and decompiled Android apps. They successfully identified 56,682 AI apps from a total of 7,259,232222The latest Android version is considered while multiple versions are stored. apps on AndroZoo. In our experiments, we leverage this dataset of 56,682 AI applications derived via the AI Discriminator approach to conduct cross-validation of our method, LLMAID.
3.1.2. Dataset 2
@@
To curate our dataset, we leverage AndroZoo and extract the latest release of each Android app using metadata from its associated CSV file. For each app, we utilize the google-play-scraper npm package to verify its current availability on Google Play Store by querying with the package name (playscraper:2023), and collect its key metadata, particularly the app description and release date. In total, we identified 2,298,423 actively maintained Android apps out of a total of 23,321,823 apps available in AndroZoo at the time of our study. To focus our evaluation on apps with likely AI functionality, we further refined the dataset by examining the app descriptions retrieved from Google Play, selecting only those that strongly indicated the presence of AI-related features. Dataset 2 complements Dataset 1 by capturing newer Android apps with potential AI functionalities beyond rule-based detection, enabling a more rigorous and forward-looking assessment of LLMAID in an open-world setting.
Existing studies have often relied on manual inspection to identify AI-enabled applications; however, such an approach is inherently time-consuming and difficult to scale to the size of popular app repositories. In practice, publicly available Android applications typically include short descriptions that summarize their core functionalities, offering a lightweight entry point for automated analysis. For example, the description of ChatGO - AI Chatbot Assistant [^3] https://play.google.com/store/apps/details?id=com.chatgpt.aichat.gpt3.aichatbot mentions that ‘Discover ChatGO - AI Chatbot Assistant, the optimized [AI assistant] that transforms the way you work, create, and connect! Built on cutting-edge technologies … writing, designing, summarizing, or just exploring new ideas, ChatGO is your [go-to AI companion] … and highly personalized assistance’. This example illustrates that AI-related applications include explicit references to AI-related terms (as highlighted in red) within their descriptions. Motivated by this observation, we design a two-stage process for identifying AI applications. In the first stage, a rule-based filter is applied to detect potential AI apps based on the presence of AI-related keywords in their descriptions. In the second stage, we employ ChatGPT to perform a deeper semantic analysis of the filtered descriptions, thereby verifying whether the applications are likely to incorporate AI technologies beyond superficial references to AI-related keywords.
Specifically, the first filtering stage applies rule-based keyword extraction to identify potential AI applications based on their descriptions. The keyword set covers a broad spectrum, ranging from high-level terms (e.g., “AI” and “Deep Learning”) to more task-specific ones (e.g., “text classification” and “facial recognition”). By performing exact matches against this vocabulary, the filter effectively narrows the candidate pool, enabling subsequent analysis to concentrate on apps with higher potential relevance. However, the mere occurrence of such keywords does not necessarily imply that AI constitutes a core functionality of the app. To address this limitation, we introduce a second filtering stage powered by ChatGPT, which interprets the semantic context of app descriptions to assess whether AI techniques are likely to be employed. This stage improves the precision of the identification process by moving beyond simple keyword occurrence matching. Through this two-stage design, our approach mitigates the limitations of purely rule-based methods and more accurately detects AI-enabled Android applications. Nonetheless, due to semantic inflation and the frequent misuse of AI-related terms (e.g., “deep learning”), even when combining keyword matching with semantic analysis of descriptions, reliance on literal text alone remains inadequate and can lead to inaccurate identification. Therefore, this dataset is treated as an open-world setting and is used to evaluate the proposed approach, where incorporating implementation details enables more reliable detection of true AI apps.
Figure 4. The distribution of collected apps over the years.
In total, we successfully identified 15,517 Android applications (considering only the latest released version) from a corpus of 23,321,823 apps available on AndroZoo. Furthermore, we extracted release dates from the metadata provided on the Google Play Store and analyzed the temporal distribution of these apps, as shown in Fig. 4. The results indicate that the number of released apps has grown in parallel with the increasing prevalence of AI technologies over the past decade. Notably, Fig. 4 presents only 15,055 applications, as release date metadata was unavailable for 462 apps, accounting for 2.98% (462 out of 15,517) of the total. This minor omission does not affect the overall observation, and the upward trend of AI technology adoption in mobile applications over the past decade remains evident.
3.2. RQ1: To what extent can LLMAID reliably and accurately determine AI capabilities in previously identified AI Apps?
In this research question, we examine whether LLMAID can effectively identify AI capabilities within a set of pre-labeled AI applications. For this purpose, we employ Dataset 1 as our evaluation benchmark. Since Dataset 1 identifies AI apps but does not provide ground-truth labels for their underlying AI components, we randomly sample a subset of 382 applications and conduct human validation to assess the accuracy of LLMAID. The sample size is calculated based on the popular online Sample Size Calculator [^4] https://www.surveysystem.com/sscalc.htm with a confidence level of 95% and a margin of error of 5%.
We first present the quantity distribution of three representative types of components, including packages, APIs, and HTTPS requests in Fig. 5. Specifically, Fig. 5 shows the total number of packages, APIs, and HTTPS requests after completing Task 1 (Candidate Extraction) of LLMAID, while Fig. 5 presents the identified AI-related packages, APIs, and HTTPS requests after finishing Task 3 (AI capability analysis and detection) of LLMAID. The median values for the total considered packages, APIs, and HTTPS requests in Fig. 5 are 433.5, 296, and 53.5, respectively. It is also worth noting that the considered packages, APIs, and HTTPS requests are proactively filtered results obtained by applying a self-curated whitelist of commonly observed non-AI components during the candidate extraction process (i.e., Task 1). The whitelist of non-AI components comprises the official Java and Android kernel-related packages and API signatures. After performing AI capability analysis and determination, the median number of AI-related packages, APIs, and HTTPS requests drops to 9.5, 33, and 0, respectively. Notably, packages are filtered more aggressively than APIs, which is expected given that many AI-specific APIs are concentrated within a smaller set of AI-related packages. In contrast, HTTPS requests are sparsely detected, primarily due to the relatively recent adoption of AI-related network calls and their limited presence in the collected dataset.
@@
Figure 5. The determination of AI candidates.
@@
To evaluate the reliability of LLMAID in identifying AI components within the sampled dataset, we conducted a manual annotation study involving two independent authors, who independently labeled a total of 49,371 components extracted from the 382 sampled AI apps. When disagreements arise, dedicated meetings with an additional author are held to facilitate consensus. Each component was classified as either an AI component or a non-AI component. We defined a True Positive as a component correctly identified by LLMAID as AI-related, while a True Negative referred to a component correctly identified as non-AI. Following human validation, LLMAID achieved a precision of 98.05% and a recall of 93.31%, indicating strong performance in correctly detecting AI components. Furthermore, the inter-annotator agreement, measured using Cohen’s unweighted kappa coefficient (cohenkappa), was 0.82, indicating substantial consistency among the human evaluators. These results collectively demonstrate the high accuracy and robustness of LLMAID in detecting AI components.
@@
In addition, we validated the reliability of LLMAID in identifying AI apps. We defined an app as an AI app if it contains one or more AI components determined by LLMAID. Our results showed that LLMAID correctly identified 378 out of 382 sampled applications as AI apps, corresponding to an accuracy of 99.0% (378/382). For the remaining four applications, we further manually inspected them and observed that the only potential AI indicators were filenames with the .model suffix. Without further complementary information, such as associated APIs, packages, or functional descriptions, LLMAID was unable to classify these apps as AI apps, hence excluding them from the set of AI apps.
@@
Table 1. Comparison of the reliability of LLMAID between the implementations of analysis-then-detection and detection-then-analysis.
Analysis-then-Detection Detection-then-Analysis
@@
Although the design of Task 3.1 (AI Capability Analysis) and Task 3.2 (AI Component Determination) is inspired by Chain-of-Thought prompt (wei2022chain), this sequence may not be optimal for LLMAID. A more efficient alternative is to first determine whether a component is AI-related or not and then analyze and interpret only those classified as AI components, effectively reversing the order of Tasks 3.1 and 3.2. Under the same evaluation setup on the 382 sampled apps, we find that this detection-then-analysis approach reduces the average processing time per app by 65%, as summarized in Table 1. However, the number of AI components identified using this approach is notably lower than that obtained with the original sequential implementation (Analysis-then-Detection). Based on the overall results, we suggest that detection-then-analysis can only be used when developers are interested in finding AI apps. If developers seek to understand all possible AI capabilities of an app, we foresee that the original LLMAID with analysis-then-detection is more effective in capturing AI candidates.
@@
Table 2. Number of AI and Non-AI apps classified by AI Discriminator and LLMAID.
@@
AI Non-AI AI Non-AI
555 3264 1366 2453
Table 3. Detected AI Apps by AI Discriminator and LLMAID
@@
3.3. RQ2: How does LLMAID compare to a state-of-the-art rule-based approach in reliably identifying AI capabilities within in-wild Android Apps?
In this research question, we evaluate how reliable LLMAID is in determining AI capabilities and compare it to the SOTA approach, AI Discriminator (li2022ai; aiapp:2023). To efficiently perform the evaluation, we select a representative subset of Dataset 2, focusing on the apps published in the recent five years, and randomly sampling 30% apps in each year to preserve the same quantity distribution. Finally, we ended up with 3,819 $(1980\times 30\%+2336\times 30\%+2326\times 30\%+3053\times 30\%+3044\times 30\%)$ apps for the experiment.
Following the same experimental setup as in RQ1, we compute the precision and recall for AI component identification. LLMAID achieves a precision of 97.32% and a recall of 91.01%, demonstrating strong performance in identifying AI components. The inter-annotator reliability, measured using Cohen’s unweighted kappa coefficient, is 0.80, which reflects a high level of agreement with the results generated by LLMAID.
Table 3.2 presents the experimental results for AI App identification, comparing the performance of AI Discriminator with that of LLMAID. Among the total of 3,819 experimental Android Apps, 555 of them are classified by AI Discriminator as AI Apps, while 1,366 are recognized as AI Apps by LLMAID, which is 242% more than AI Discriminator. Table 3.2 shows the number of AI Apps identified by AI Discriminator and LLMAID. It clearly demonstrates that 546 Android Apps, accounting for 90.55% $(546/603)$ of those identified by AI Discriminator and 39.97% $(546/1366)$ of those confirmed by LLMAID, are commonly recognized as AI Apps by both approaches. To manually inspect the nine AI apps identified solely by AI Discriminator, multiple authors downloaded the APK files from AndroZoo and dissected them with Jadx (jadx2023). Two apps are recognized as AI apps because only one special tflite file is being detected, while the remaining seven apps are classified as AI apps because of keywords, such as “DecisionTreeClassifier” and “KNeighborsClassifier” in their descriptions of json files. However, the description contents in the JSON files alone, without supporting code-level detection, may not imply the actual implementation of AI capabilities and therefore should not be regarded as convincing AI evidence.
To verify the 820 apps only classified by LLMAID, we randomly sampled 86 apps (with a confidence level of 95% and a margin of error of 10%) and cross-validated them using the decompiler tool, Jadx (jadx2023), with multiple authors involved. In the course of this cross-validation, we observed that OpenAI APIs were sometimes directly invoked within Java code, whereas in other cases developers implemented these API calls through native libraries (i.e.,.so files). Such implementations of cloud-based AI capabilities could not be detected by AI Discriminator due to the limitations of its rule set. In the end, we confirmed that 83 of the total of 86 sampled apps are classified by LLMAID as truly AI apps, showcasing a high level of identification precision.
@@
Figure 6. Description of Copy.ai (the left) and its AI service summarization (the right) generated by our framework.
3.4. RQ3: How is the usefulness of the AI service summary generated by LLMAID?
In this research question, we seek to evaluate the usefulness of the AI summary, which is generated in Task 4 of LLMAID based on the app’s implementation details. Specifically, we compare the generated AI summaries with the developer-provided descriptions on the app store, using this comparison as a diagnostic tool to reveal the shortcomings of the developer-provided information and to enable a more accurate understanding and evaluation of AI functionalities.
We conducted a user study by recruiting six developers, each with at least three years of experience in Android app analysis. Specifically, we randomly selected 20 Android apps identified as AI apps in RQ2, and then generated their AI summaries. For each app, we paired each LLMAID generated AI summary with its corresponding app description retrieved from the Google Play store via the google-play-scraper (playscraper:2023), and displayed them side-by-side to participants, as shown in Fig. 6. We divided the six participants as two groups. Each group evaluated ten Android apps and each app was evaluated by three participants individually. Before conducting the user study, participants were instructed to download the apps and get familiar with their major features. We then asked participants to read and compare the app description and the AI summary for each app, evaluating both from five aspects: Accuracy (i.e., Do you think that the summary is accurate concerning the AI functions the app provides?), Completeness (i.e., Do you think that the summary is complete about the provided AI tasks?), Consistency (i.e., Do you think that the summary is concrete enough concerning the AI functions?), and Recommendations for app users (i.e., Would you like to recommend the summary to other app users?) and developers (i.e., Would you like to recommend the summary to other app developers to inspire new app development?). Each dimension was assessed using a 10-point Likert scale, with scores ranging from 1 (indicating poor quality) to 10 (indicating high quality).
Table 4. Average and median scale values voted for sampled Android apps by six participants. “Recom.” stands for Recommendation, “devs” stands for developers, “Med.” stands for median.

### v2

@@ 3 Evaluation
In this section, we evaluate the effectiveness of our approach, LLMAID, by addressing the following research questions:
@@
    This research question aims to assess the accuracy and reliability of LLMAID in analyzing AI capabilities within previously recognized AI Android apps. Addressing this question validates the core effectiveness of LLMAID and establishes a foundation for further evaluation.
@@
    We explore the clarity and informativeness of the AI service summary provided by our framework. This summary provides end users, developers, and regulators with an insightful overview of the AI functionalities of a given Android app.
@@ 3.1 Dataset
We perform our experiments on two separate datasets. The first one comes from the existing literature, which was collected from AndroZoo and contains 56,682 rule-based approach determined AI Android apps allix2016androzoo. The other one is self-curated, comprising 15,517 Android apps, to address the absence of recently released applications, especially those integrated with emerging LLMs such as ChatGPT. We present these two datasets in the following paragraphs.
3.1.1 Dataset 1
@@
Li et al. li2022ai proposed a SOTA rule-based approach, AI Discriminator, to identify AI-driven Android apps via keywords matching on the dissected and decompiled Android apps. They successfully identified 56,682 AI apps from a total of 7,259,232222The latest Android version is considered while multiple versions are stored. apps on AndroZoo. In our experiments, we utilize the dataset of 56,682 Android AI apps to perform cross-validation of our approach.
3.1.2 Dataset 2
@@
To curate our dataset, we leverage AndroZoo and extract the latest release of each Android app using metadata from its associated CSV file. For each app, we utilize the google-play-scraper npm package to verify its current availability on Google Play Store by querying with the package name playscraper:2023, and collect its key metadata, particularly the app description and release date. In total, we identified 2,298,423 actively maintained Android apps out of a total of 23,321,823 apps available in AndroZoo at the time of our study. To focus our evaluation on apps with likely AI functionality, we further refined the dataset by examining the app descriptions retrieved from Google Play, selecting only those that strongly indicated the presence of AI-related features.
Previous studies often rely on manual inspection to identify AI apps, however, this approach tends to be time-consuming and challenging to scale. A publicly available Android app conventionally includes a brief description that clarifies its core functionalities. For example, the description of Wise AI Chatbot [^3] https://play.google.com/store/apps/details?id=com.jgapps.chatgpt mentions that ‘Wise Chat GPT Pro is the first app to utilize … ChatGPT 3.5 Turbo to provide instant and [intelligent] responses … with Wise Chat GPT Pro, our [AI] [chatbot] powered by ChatGPT 3.5 Turbo technology …’. This example illustrates how AI-related apps often explicitly reference AI-related terms (i.e., highlighted in red) in their descriptions. Motivated by this observation, we transform this step into a two-staged process. We begin by applying a rule-based filter to identify potential AI apps through the presence of AI-related keywords in their descriptions. Subsequently, we employ ChatGPT to perform a deeper semantic analysis of these filtered descriptions, further verifying whether the apps likely incorporate AI technologies.
To be specific, the filtering stage employs rule-based keyword extraction to identify potential AI apps based on their application descriptions. We define a wide range of keywords from high-level functionalities (e.g., “AI” and “Deep Learning”) to low-level ones (e.g., “text classification” and “facial recognition”). By matching these terms exactly, the filter effectively reduces the candidate pool, allowing subsequent analysis to focus on high-potential AI apps. However, the mere presence of such keywords does not always indicate that AI constitutes a core functionality of the app. To address this limitation, we introduce a second filtering stage using ChatGPT, which evaluates whether the app is likely to employ AI techniques based on its description. This stage further enhances the precision of our identification process by interpreting the context of the apps’ descriptions. By employing the two-stage process, our approach overcomes the limitation of keyword matching methods, hence accurately detecting AI Android apps.
Figure 4: The distribution of collected apps over the years.
@@
Figure 5: The determination of AI candidates.
In total, we successfully identified 15,517 Android apps (only the latest released version is considered) from a total of 23,321,823 apps on AndroZoo. Additionally, we extracted the release date from the metadata provided on the Google Play Store and presented the distribution of these apps over the last decade in Fig. 5. Our data shows that the growth in released apps aligns closely with the increasing prevalence of AI technologies during this period. Notably, Fig. 5 includes only 15,055 apps, as release date metadata was unavailable for 462 apps, representing just 2.98% (462 out of 15,517) of the total. Despite this, the upward trend of employing AI technologies in mobile apps over the past ten years remains valid.
3.2 RQ1: To what extent can LLMAID reliably and accurately determine AI capabilities in previously identified AI Apps?
In this research question, we examine whether LLMAID can effectively identify AI capabilities within a set of pre-labeled AI applications. We employ Dataset 1 as our evaluation benchmark. However, since Dataset 1 only determines AI apps but does not provide ground-truth labels for their detailed underlying AI components, we randomly sample a subset of 382 AI apps and perform human validation to assess LLMAID. The sample size is calculated based on the popular online Sample Size Calculator  [^4] https://www.surveysystem.com/sscalc.htm with a confidence level of 95% and a margin of error of 5%.
We first present the quantity distribution of three representative types of components, packages, APIs, and HTTPS requests in Fig. 5. Specifically, Fig. 5 shows the total number of packages, APIs, and HTTPS requests after completing Task 1 (Candidate Extraction) of LLMAID, while Fig. 5 presents the identified AI-related packages, APIs, and HTTPS requests after finishing Task 3 (AI capability analysis and detection) of LLMAID. The median values for the total considered packages, APIs, and HTTPS requests in Fig. 5 are 433.5, 296, and 53.5, respectively. It is also worth noting that the considered packages, APIs, and HTTPS requests are proactively filtered results obtained by applying a self-curated whitelist of commonly observed non-AI components during the candidate extraction process (i.e., Task 1). The whitelist of non-AI components comprises the official Java and Android kernel-related packages and API signatures. After performing AI capability analysis, the median number of AI-related packages, APIs, and HTTPS requests drops to 9.5, 33, and 0, respectively. Notably, packages are filtered more aggressively than APIs, which is expected given that many AI-specific APIs are concentrated within a smaller set of AI-related packages. In contrast, HTTPS requests are sparsely detected, primarily due to the relatively recent adoption of AI-related network calls and their limited presence in the collected dataset.
@@
To evaluate the reliability of LLMAID in identifying AI components within the sampled dataset, we conducted a manual annotation study involving two independent authors, who independently labeled a total of 49,371 components extracted from the 382 sampled AI apps. When disagreements arise, dedicated meetings involving an additional author are held to facilitate consensus. Each component was classified as either an AI component or a non-AI component. We defined a True Positive as a component correctly identified by LLMAID as AI-related, while a True Negative referred to a component correctly identified as non-AI. Following human validation, LLMAID achieved a precision of 98.05% and a recall of 93.31%, indicating strong performance in correctly detecting AI components. Furthermore, inter-annotator agreement measured using Cohen’s unweighted kappa coefficient cohenkappa was 0.82, reflecting substantial consistency among human evaluators. These results collectively demonstrate the high accuracy and robustness of LLMAID in detecting AI components.
@@
In addition, we validated the reliability of LLMAID in identifying AI apps. We defined an app is considered an AI app if it contains one or more AI components determined by LLMAID. Our results showed that LLMAID can detect 378 out of 382 sample apps as AI apps, indicating an accuracy of 99.0% (372/382) is achieved. We further analyzed the remaining four apps and observed that the only potential AI indicators were filenames with the .model suffix. Without further complementary information, such as associated APIs, packages, or functional descriptions, LLMAID was unable to classify these apps as AI apps, hence excluding them from the set of AI apps.
@@
Table 1: Comparison of the reliability of LLMAID between the implementations of analysis-then-detection and detection-then-analysis.
Analysis/Detection Detection/Analysis
@@
Although the design of Task 3.1 (AI Capability Analysis) and Task 3.2 (AI Component Determination) is inspired by Chain-of-Thought prompt wei2022chain, it may not be optimal for LLMAID. A more efficient alternative is to first classify the components and then analyze only those identified as AI components, reversing the order of Tasks 3.1 and 3.2. Following the same evaluation setup on the 382 sampled apps, we find that this detection-then-analysis approach reduces the processing time per app by 65%, as shown in Table 1. The number of identified AI components is significantly lower than that of the original sequential implementation. Based on the overall results, we suggest that detection-then-analysis can only be used when developers are interested in finding AI apps. If developers seek to understand all possible AI capabilities of an app, we foresee that the original LLMAID with analysis-then-analysis is more effective in capturing AI candidates.
@@
3.3 RQ2: How does LLMAID compare to a state-of-the-art rule-based approach in reliably identifying AI capabilities within in-wild Android Apps?
In this research question, we evaluate how reliable LLMAID is in determining AI capabilities and compare it to the SOTA approach, AI Discriminator li2022ai; aiapp:2023. To efficiently perform the evaluation, we select a representative subset of Dataset 2, focusing on the apps published in the recent five years, and randomly sampling 30% apps in each year to preserve the same quantity distribution. Finally, we ended up with 3,819 $(1980\times 30\%+2336\times 30\%+2326\times 30\%+3053\times 30\%+3044\times 30\%)$ apps for the experiment.
Following the same experimental setup as in RQ1, we compute the precision and recall for AI component identification, and LLMAID achieves a precision of 97.32% and a recall of 91.01% in identifying AI components. The inter-annotator reliability, measured by Cohen’s unweighted kappa coefficient, is 0.80, indicating a high level of agreement with the results generated by LLMAID.
Table 2: Number of AI and Non-AI apps classified by AI Discriminator and LLMAID.
@@
AI Apps Non-AI Apps AI Apps Non-AI Apps
603 3216 1366 2453
Table 3: Detected AI Apps by AI Discriminator and LLMAID
@@
Table 3.3 presents the experimental results of AI App identification, comparing the performance of AI Discriminator and LLMAID. Among the total of 3,819 experimental Android Apps, 603 of them are classified by AI Discriminator as AI Apps, while 1,366 are recognized as AI Apps by LLMAID, which is 242% more than AI Discriminator. Table 3.3 shows the number of AI Apps identified by AI Discriminator and LLMAID. It clearly demonstrates that 546 Android Apps, accounting for 90.55% $(546/603)$ of those identified by AI Discriminator and 39.97% $(546/1366)$ of those confirmed by LLMAID, are commonly recognized as AI Apps by both approaches. To manually inspect the nine AI apps identified solely by AI Discriminator, multiple authors downloaded the APK files from AndroZoo and dissected them with Jadx jadx2023. Two apps are recognized as AI apps because only one special tflite file is being detected, while the remaining seven apps are classified as AI apps because of keywords, such as “DecisionTreeClassifier” and “KNeighborsClassifier” in their descriptions of json files. However, the description contents in the JSON files may not imply the implementation of AI capability and therefore should not be counted as convincing AI evidence.
To verify the 820 apps only classified by LLMAID, we randomly sample 86 apps (with a confidence level of 95% and a margin of error of 10%) and cross-validated by multiple authors with the decompiler tool, Jadx jadx2023. During cross-validation, we observe that OpenAI APIs are sometimes directly invoked within Java code, whereas other developers implement these API invocations via native libraries (i.e.,.so files). This API invocation implementation of the cloud AI capabilities could not be detected by the AI discriminator, given its limited rule set. In the end, we confirm that 83 of the total of 86 sampled apps are classified by LLMAID as AI apps, showcasing a high level of identification precision.
@@
3.4 RQ3: How is the usefulness of the AI service summary generated by LLMAID?
Figure 6: Description of Copy.ai (the left) and its AI service summarization (the right) generated by our framework.
In this research question, we seek to evaluate the usefulness of the AI summary, which is generated in Task 4 of LLMAID. We conducted a user study by recruiting six developers, each with at least three years of experience in Android app analysis. Specifically, we randomly selected 20 Android apps identified as AI apps in RQ2, and then generated their AI summaries. For each app, we paired each LLMAID generated AI summary with its corresponding app description retrieved from the Google Play store via the google-play-scraper playscraper:2023, and displayed them side-by-side to participants, as shown in Fig. 6. We divided the six participants as two groups. Each group evaluated ten Android apps and each app was evaluated by three participants individually. Before conducting the user study, participants were instructed to download the apps and get familiar with their major features. We then asked participants to read and compare the app description and the AI summary for each app, evaluating both from five aspects: Accuracy (i.e., Do you think that the summary is accurate concerning the AI functions the app provides?), Completeness (i.e., Do you think that the summary is complete about the provided AI tasks?), Consistency (i.e., Do you think that the summary is concrete enough concerning the AI functions?), and Recommendations for app users (i.e., Would you like to recommend the summary to other app users?) and developers (i.e., Would you like to recommend the summary to other app developers to inspire new app development?). Each dimension was assessed using a 10-point Likert scale, with scores ranging from 1 (indicating poor quality) to 10 (indicating high quality).
Table 4: Average and median scale values voted for sampled Android apps by six participants. “Recom.” stands for Recommendation, “devs” stands for developers, “Med.” stands for median.